OpenStudy (anonymous):

On Mars, an object weighs 38% as much as on Earth. How much would a person who weighs 150 pounds on Earth weigh on Mars?

OpenStudy (anonymous):

wait was tht right hero

hero (hero):

No it wasn't

hero (hero):

207 is correct. Don't ask me how I got it

hero (hero):

There are several ways to calculate percent

OpenStudy (anonymous):

oh i kinda need to know how u got it

hero (hero):

The question says that someone on Mars weighs 38% more than what someone weighs on Earth. In order to answer the question, you need to know what 38% of 150 is. In other words, some number is 38% of 150. So you use this general formula: is/of = %/100 Substituting the values, we get: x/150 = 38/100 So now, you cross multiply and solve for x as I've done above. x = 57 So if 57 is 38% of 150, we have to add that to 150 because we know that the person on Mars weighs 38% MORE. So 150 + 57 = 207. Therefore, the person on Mars weighs 207 lbs.

OpenStudy (lilg132):

theres an easier way i think